OpenStudy (anonymous):

Solve the equation??? (: 0.5x-1.4=1.1

OpenStudy (jonnyvonny):

I'm assuming you wish to know when the function will be zero. To do this, you have to get "x" alone. When getting a variable alone, you must follow the reverse of PEMDAS. Thereby, you first must add "1.4" to both sides to get ".5x=1.1". Lastly, you divide both sides by ".5". The final answer is "5". If you need more elaboration, feel free to message. :)

OpenStudy (anonymous):

thanks you teehee (:

OpenStudy (anonymous):

0.5x - 1.4 = 1.1 (add 1.4 to each side) 0.5x = 1.1 + 1.4 0.5x = 2.5 (divide both sides by 0.5) x = 2.5/0.5 x = 5 I agree :)